OpenStudy (nuccioreggie):

What is the measure of angle x? A right angle is shown divided into two parts. The measure of one part of the right angle is 30 degrees. The measure of the other part is 3x. 10 20 30 40

OpenStudy (nuccioreggie):

@welshfella

OpenStudy (nuccioreggie):

@triciaal

OpenStudy (anonymous):

Start with the measure of a right angle. Do you know what that is? When you know that, then you know there are 2 parts, the first is 30, the other is unknown, together they add to the measure of a right angle. So your answer would be the measure of a right angle - 30

OpenStudy (anonymous):

Sorry, I missed the 3x part. So your answer would be 3x = right angle - 30, solve for x.

OpenStudy (nuccioreggie):

So its C Correct?

OpenStudy (anonymous):

No. The answer is not c. Show your work. How did you get to C as your answer?

OpenStudy (nuccioreggie):

No its A because i divided the two number and got 10

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (anonymous):

The answer is not A either. What two numbers did you divide?

OpenStudy (nuccioreggie):

30 and 3 so that means that the anser is B

OpenStudy (nuccioreggie):

@psychotink

OpenStudy (anonymous):

You do not divide 30 by 3. You missed a part. You first find the measure of a right angle. Really what you are looking at here is: Right angle=30 + 3x Plug in the measure of a right angle and solve.

OpenStudy (nuccioreggie):

Ok so D

OpenStudy (nuccioreggie):

@Psychotink

OpenStudy (anonymous):

Are you trying or just guessing? Please post exactly what you are doing to get your answer.

OpenStudy (nuccioreggie):

No u are wrong about the stuff u said was incorrect u said it wasnt B but it actually is B i did the math got got this. it is 3x=90-30 3x=60 x=20 N its B so i think u should re check ur work before giving false information hahahahahahahaha

OpenStudy (anonymous):

I never said it wasn't B. I said you don't divide 30 by 3. That was never false information. But I wasn't going to give you the answer without you knowing how it happened first.